Ayayai Inc. manufactures a variety of consumer products. The company’s founders have run the company for 30 years and are now interested in retiring. Consequently, they are seeking a purchaser who will continue its operations, and a group of investors, Morgan Inc., is looking into the acquisition of Ayayai. To evaluate its financial stability and operating efficiency, Ayayai was requested to provide the latest financial statements and selected financial ratios. Summary information provided by Ayayai is as follows.

Ayayai Inc.
Income Statement
For the Year Ended November 30, 2021
(in thousands)
Sales (net) $ 30,590
Interest income 590
      Total revenue 31,180
Costs and expenses
  Cost of goods sold 17,690
  Selling and administrative expenses 3,640
  Depreciation and amortization expense 1,980
  Interest expense 990
      Total costs and expenses 24,300
  Income before taxes 6,880
    Income taxes 2,752
  Net income $ 4,128
Ayayai Inc.
Balance Sheet
As of November 30
(in thousands)
2021 2020
Cash $ 490 $ 590
Short-term investments (at cost) 390 290
Accounts receivable (net) 3,290 2,990
Inventory 6,090 5,490
    Total current assets 10,260 9,360
Property, plant, & equipment (net) 7,190 7,090
    Total assets $ 17,450 $ 16,450
Accounts payable $ 3,790 $ 3,490
Income taxes payable 990 890
Accrued expenses 1,790 1,490
    Total current liabilities 6,570 5,870
Long-term debt 2,090 1,890
    Total liabilities 8,660 7,760
Common stock ($1 par value) 2,790 2,790
Paid-in capital in excess of par 1,090 1,090
Retained earnings 4,910 4,810
    Total stockholders’ equity 8,790 8,690
    Total liabilities and stockholders’ equity $ 17,450 $ 16,450
Selected Financial Ratios
Ayayai Inc. Current
Industry
Average
2020 2019
Current ratio 1.59 1.60 1.72
Acid-test ratio 0.66 0.59 0.82
Times interest earned 8.64 8.59 8.54
Profit margin on sales 13.29 % 12.19 % 13.09 %
Asset turnover 1.93 1.92 1.93
Inventory turnover 3.26 3.3 3.27


(a)

Calculate a new set of ratios for the fiscal year 2021 for Ayayai based on the financial statements presented. (Round answers to 2 decimal places, e.g. 52.75 or 52.75%.)

Current ratio : 1
Acid-test ratio : 1
Times interest earned
Profit margin on sales %
Asset turnover times
Inventory turnover times
Answer #1

Current Ratio = Current Assets / Current Liabilities
= $10260 / 6570 = 1.56 : 1

Acid-Test Ratio = (Current Assets - Inventory) / Current Liabilities
= ($10260 - 6090) / 6570 = 0.63 : 1

Times Interest Earned = (Income before taxes + Interest Expense) / Interest Expense
= ($6880+990) / 990 = 7.95 times

Profit Margin on Sales = Net Income / Sales
= $4128 / 30590 = 13.49%

Asset Turnover = Sales / Average Total Assets
= $30590 / 16950 = 1.80 times

Inventory Turnover = Cost of Goods Sold / Average Inventory
= $17690 / 5790 = 3.06 times

If you have any query, kindly comment with your query and please mark thumbs up.